Fermi National Laboratory
 

Fermilab Shutdown Report for July 17, 2009 - July 24, 2009

 

Fermilab Shutdown Report

7/17/09 – 7/24/09

LINAC

Shutdown Week 6 – On Schedule

  • Preaccelerator
    • Both dome walls were cleaned
    • Progress with the new pulse shifter box
    • Generator installed in the H- Source dome
    • New source installed in the H- system.
  • Low Energy
    • Cleaned/Regreased the PA filament inductrol cabinet
    • Performed electrical tests on stations L RF4 & LRF5
    • PA changed out on LRF1 & LRF2
  • High Energy
    • Charging supply and PFN hi pot after maintenance
    • Problems with Klystron (KRF6) PFN
    • Working with people to mitigate the situation
  • Tunnel Activities
    • Water leak tests on DT20Tk5
    • Stem box closed on tank 5
    • Replaced TK1COL read back device
    • Terminate HV BLM cables
    • Cleaning crew cleaned the tunnel
    • Safety system tests performed for Linac/NTF
  • Next Week:
    • Low Energy start up after the outage in preparation of initiating Neutron Therapy Facility (NTF) treatment (starting week August 3 rd)
      • (beam up to NTF will be requested next week for tuning)
    • Continue with HE activities

Booster

Shutdown Weeks 6 – On Schedule

  • Correction element upgrade installation 100% complete
    • 36 out of 36 magnets installed and aligned
    • Vacuum pump down is progressing, although some areas are slow to improve
    • Corrector wire pulls 100% complete
  • East Gallery LCW Cooling pipe upgrade 100% completed. Need to hydro-test.
  • GMPS work 85% complete.
  • L12 kicker magnets installed.
  • Laser Profile Monitor installed
  • Corrector magnet work continues with electrical and water connections. Database updates progressing.
  • Vacuum pump down, leak checking.
  • CUB filling the Booster LCW system and have supply pumps running to circulate water.
  • FESS replacing the sump line isolation valve to the Booster pond.
  • Monday tunnel cleaning crew is scheduled

Antiproton Source

Accumulator 4-8GHz DP Kicker Upgrade: Ø~95% complete

The A50-A60 bake is done. There remain some vacuum system controls issues and reconnection of stochastic cooling electronics

Debuncher Septum Magnet Replacement: Ø~70% complete

The broken magnet has been repaired and has passed a hi-pot. Today it will be lowered back into the tunnel and transported to the D40 area for installation

IB7 (AP2 dipole magnet) Replacement: Ø98% complete

The new magnet is in place along with our repaired momentum collimator. Survey is complete and most of the AP2 line should be back under vacuum by the end of today

Other Work

  • Maintenance work on various RF systems continues
  • LCW leaks being addressed. Re-brazing of sextupole magnet leaks completed
  • Target SEM work continues
  • New door installed in U/S AP1 enclosure

Main Injector

MI-10 area: On Schedule

  • All conduits in place, grouting continues, backfilling in progress (not quite to 5ft)
  • Building foundation progressing quickly

MI-40 area: On Schedule

  • All 12 conduits in place
  • Berm partially restored
  • Building foundation started

Tunnel Work

MI-8

  • Beam valves relocated
  • Multiwires being assembled/tested/prepped for installation
  • New ceramic MW @ 810, 811, 813, 814 & 830 for E + dp/p measurements. (maybe 851)
  • Two Texas multiwires being replaced due to vacuum leaks

MI-10

  • Tunnel closed until later today?

MI-40

  • Beam valve and kickers moved
  • Gap Clearing kickers being assembled
  • Kicker stand installation continues
  • Thinking of adding 2 more BPMs in this area for eCloud

MI LCW

  • 12 quad locations left to re-hose
  • Aligning/repairing LCW pumps around the ring
  • Replacing/calibrating pots in 3-way valves
  • RF heat exchangers cleaned

MI PS

  • Too much work to list. All maintenance is going well: on schedule

MI eCloud (MI-52)

  • Added eCloud BPMs at 518, 519
  • Beam valve installed @522
  • Preparing to install eCloud components

MI RF

  • Modulators cleaned, strainers cleaned, all work on schedule

New Work

  • Fixed vacuum leak (609-619) caused by non MI work in the area

Recycler

Recycler

  • Pot RR129 (i.e. MI-10 injection kicker)
    • Prep work started
      • Got delayed a bit, Complete by Wednesday?
  • Move 6 magnets ( complete)
  • Vacuum maintenance
    • Leak checking 800 line
      • No air leaks found so far but RGA scan shows high Argon peak
  • Power supplies maintenance
    • In progress
  • CIA crate activities (80% complete)
  • HLS installation at MI-10
    • ALARA plan has been reviewed
    • Plan on installing hardware when construction work in the tunnel is completed

Electron Cooling

  • Pelletron maintenance (50% complete)
    • All mechanical work inside the tank but the 4” GVM installation has been completed
      • Chain running test successfully completed on Tuesday
      • Collector water system purged, cleaned and refilled
    • Electrical work mostly completed
      • Minor fixes
  • Flush Water/Glycol skid
    (and magnets) (90% complete)
    • Transfer skid water loop has been isolated from the cooling skid water loop
    • Finish filling both systems with water/glycol mix today
  • NMR probes replacement (90% complete)
    • One last probe needs to be tuned in
  • Transferring gas on August 19 th
    • 2-3 days of conditioning
    • 4-5 days U-Bend mode
      • Check protection system
      • Check controls, regulation loops…
      • Miscellaneous measurements
  • Then, more to do when MI tunnel is secure

Recycler Pump Out

  • Pumping cart moved from E4R to MI-40 on Tuesday
    • Manifold hooked to 405 roughing port and baked overnight
  • Unexpected drilling work on Wednesday
  • RGA scans made on Wednesday afternoon and Thursday morning turned out to be different
    • Some water seems to be ‘leaking’ in
    • More scans and careful review by experts in progress to determine if we can go ahead with the pump out or not

Tevatron

Other Items

  • Un Roll TeV magnets 3 to complete all rolls >2 mr
  • E-3 Feed can done conning tower on E35-1 spool repaired
  • Working in F1 now
  • Vertical crystal collimator bake out nearly complete
  • Repairing HLS systems around Tevatron D0 system up B0 partially, complete need to have IMU closed to finish
  • Turbo pump work around the ring
  • Sump pump and dehumidifier maintenance around the ring
  • Will have day and night shift for vacuum work starting July 27

NuMI

Shutdown Weeks 5

  • Inspect & Photograph Decay Pipe Window
  • Remove RAC Samples from Target Chase
  • Remove two (2) old stripline blocks from Target Hall and move to TSB for storage
    • Including storing stripline blocks in custom made coffins
  • Install backup water pumps & sumps in the east and west Pre-Target drains
  • Remove all R-Blocks from the Target Chase in preparation for survey of Horn 2, Horn 1, and the Target
  • EPB Dipoles PEEK Installation Continues
    • Work Continues on HV101-3, HV101-4, HV101-5
  • Install fixture for Hadron Monitor Replacement
  • Continue Installation of Controls System for Absorber Hall Dehumidification System
  • Continue installation of rope filter system for ground water cooling leg to Absorber Hall

Cryogenic System

TeV at LN2 Temperatures

Operations and Maintenance Status

  • Continued support AD/MSD for A3, D3, E3, F1 repairs
  • Reconfiguring A3 and D3 back to LN2 operations. Cooldown after 7/27 power outages. E3 feed can is replaced and cryo KVs and VPTs are reconnected
  • Continued ringwide vacuum pump maintenance, repairs, and pumpdown of various components to assure stable Tevatron LN2 Mode Operations
  • 22 engines complete; 15 flywheels replaced (A1, A2-flywheel, A3, A4, B1-flywheel, B2-flywheel, B3-flywheel, B4-flywheel, C1-flywheel, C2-flywheel, C3-flywheel, C4-flywheel, D1, D2-flywheel, D3-flywheel, D4-flywheel, E1, E2, E4, F2-flywheel, F3-flywheel, F4-flywheel, P-Bar Dry has been overhauled and P-Bar wet with new flywheel), 5/5 wet engine motors replaced
  • Mycom Compr Benshaw starter modifications all complete. Continuing 120V upgrade for Cutler-Hammer electronics
  • Mycom Compr relief valves replaced (3/8). Need to pull/replace 6 Mycoms. 4/6 IHI cold compressors replaced; one IHI has been overhauled; C4 IHI will be installed next week
  • Ongoing numerous FRIG/Compr controls/instrumentation maintenance/repairs, incl. resistors, vac TC, wiring, fans, He sniffers, etc.
  • CHL at room temperature & maintenance continuing (equip greasing/oil changes, NRL Niagara aftercooler cleaning-50% done, relief valves inspection/testing, instr. repairs.
  • Maintaining LN2 supply to TeV relying on Linde deliveries (2-3 per day) to CHL’s 150,000 gallons of tank storage
  • Completed nitrogen booster compressor maintenance
  • Scheduled recycle TSI 2.5MW compressor overhaul for Aug 10 by outside subcontractor
  • Expect delivery of repaired turbo expanders for the CHL cold box on Aug 17

Controls

Completed This Week:

  • Prep for weekend control system outage
  • Linac Controls Maintenance
  • General Controls Maintenance

Next Week:

  • •D Networks, file servers, print servers, Internet access, CAMAC FEs, Clock distribution & Java nodes will be down all weekend beginning at 5:30 Friday, 7/24/09
    • Basically, no control system until power and cooling are fully restored Monday morning
  • Control system recovery Monday morning
  • Linac startup support
  • General Controls Maintenance

MTA

The permanent magnet build has been completed.

FESS

Electrical and High Voltage

  • MI – HV & electrical PMs completed included FDRS 71/96/97 for transformers, switchgear, panels, etc.
  • Findings at MI40 pulse power equipment – very dirty – installed air filters
  • TG – HV & electrical PMs completed on FDR 43
  • Re-lamping continued on 4 -12 shift @ Pbar into next week to include SWYD

Mechanical

  • TEV - F-3 pond pump [30hp] – required packing gland replacement
  • TEV - E O –B pond pump [75hp] – installed vibration damper off of motor for improved performance
  • TEV - EO - A pond pump [75hp] – requires replacement – poor pumping performance.
  • Booster ring sump pump – requires replacement

HVAC

  • TEV – Marvair AC units PM completed
  • MI – continued dehumidifier PM
  • HVAC tech coverage on Sat & Sun for cooling SD needs

CUB

  • LCW pump PMs work in progress (WIP)
  • Chiller #3 status – Trane reviewed repair options of the 30 leaking tubes [between tube sheet] & recommends tube & tube sheet replacement. Will need to meet/discuss with FESS/AD management
  • Chiller #4 – condenser/evaporator cleaned & ready for eddy current testing
  • Chiller # 5 – cleaning WIP
  • HE-1 booster plate/frame heat exchanger – cleaned & ready to open for plate inspection
  • HE-9 Pbar plate/frame heat exchanger – cleaned
  • Booster return 95 LCW system – Bellows & 3 valves installed. The 6 valves & 1 compensator are being replaced for the middle & north older pumps. Pipefitters are working this Saturday to complete prior to the Tuesday 7/28 date.
  • Cub Shutdown starts at 6:30 am Saturday thru 8pm Sunday

Other Work and Comments

  • CASEY’S POND SHUT DOWN SATURDAY WORK 8AM TO 4PM ZEBRA MUSSEL INSPECTION - ICW valve configuration will begin at 6:00 am which means ICW via C4 pump house will only flow to FCC [STARTING AT 8AM] during this time [Casey’s pumping to be scheduled off with the re-configuration] period which means there will be virtually no cooling or fire protection to the main site with the exception of MINOS tunnel water injection into the ICW [~130gpm] in the accelerator area.
  • ICW FLUSH followed by Zebra Mussel treatment starts 8/1 thru 8/8

Switchyard 120 – Work has not started

Collider Detector at Fermilab

  • The end plugs will put in next week

D-Zero – On Schedule

Mechanical

Overall, well, but MI has been a struggle.

Instrumentation – On Schedule

  • TLMs have been installed in Linac
  • The multiwire parts will be delivered on Monday

Alignment

FESS

  • Survey of the berm for radiation shielding at MI10 after the civil construction is complete
  • Survey of the berm for radiation shielding at MI40 after the civil construction is complete

MTA

  • Align new permanent magnet to be installed on the Linac side. See experiment for details
  • Survey of the berm for radiation shielding at MI40 after the civil construction is complete

Main Injector

  • Align 7 new kicker magnets to be installed in the MI40 area. See experiment for details. 7 Components
  • After the shutdown is complete we may need to move magnets to adjust our orbits in the Main Injector and Recycler. This could be as much as 5 magnets in each machine (10 total) Shutdown 2009
  • As-found all the Main Injector and Recycler components in cells102-109. The essential components are: Recycler = magnets and cooling tanks no trims. 40 Components
  • As-found all the Main Injector and Recycler components after construction back fill in cells 341-404. Align all kickers and new beam pipe. 40 Components

MTA

  • Align new permanent magnet to be installed on the Linac side. See experiment for details
  • Survey of the berm for radiation shielding at MI40 after the civil construction is complete

NuMI

  • Align new spare Target 2 (NT-03) in target hall. 1 Component
  • Realign 6 EPB dipoles @ (HV101 magnet string). 6 Components
  • Align 3 replacement Fermi multiwires at PM101, PM105, & PM1143 Components
  • Align a new Texas multiwire at PMTGT. 1 Component
  • Re-align Fermi multiwire at PM121 after stand adjustment (and re-check adjacent instrumentation). 1 Component
  • Layout the location of 6-7 magnets in the NuMI Stub for LBNE (DUSEL) beamline magnet positioning optimization (estimated total time 1day). 7 Components

D0

  • Stickmike survey of Muon system including CF and EFs after closing detector. The start and finish dates are preliminary and depend on shutdown progress

Switch Yard

  • As found 2 EPB dipoles in enclosure B (VH94) before reworked EPB dipoles. Align 2 new dipoles

Safety

Safety

  • Shutdown ALARA plans are progressing very close to as planned
  • All MI excavation materials are back in place above the enclosure
  • All Beam Permits and Run Conditions have been rescinded
    • Accelerator restart sheets are available from ES&H

Interlock Work

  • HTC and HINS cave ODH system work
    • Completed
  • Single Line Electrical Drawing Verifications
    • Pbar 90% complete
    • Tevatron not started
  • Misc. Tevatron Interlock repairs
    • 50% complete
  • Tevatron ground fault diagnostics installation
    • 20% complete
  • Safety System Testing
    • Linac and NTF completed
    • Working to schedule the Pelletron for next week

Shutdown Safety

  • Injuries
    • Investigating an incident where two workers were sprayed with compressor oil while changing an oil filter
    • One worker oil on skin - no injury
    • The other oil in the eyes - oil was flushed out locally and then again by the medical office
      • First Aid Case

Power Outages on Monday

  • 0700-0730 power will be off to all accelerator tunnels except for MI
  • 0800-0830 power will be in the MI
  • Workers are not to be in the tunnels during the power outages

Continue using the Take Five principles before, during, and after work activities to:

  • Avoid Injuries
  • Reduce Environmental Impacts
  • http://www.fnal.gov/takefive/
 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 07/24/2009    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory